### v4

@@ Abstract
Pneumonia, a prevalent respiratory infection, remains a leading cause of morbidity and mortality worldwide, particularly among vulnerable populations. Chest X-rays serve as a primary tool for pneumonia detection; however, variations in imaging conditions and subtle visual indicators complicate consistent interpretation. Automated tools can enhance traditional methods by improving diagnostic reliability and supporting clinical decision-making. In this study, we propose a novel multi-scale transformer approach for pneumonia detection that integrates lung segmentation and classification into a unified framework. Our method introduces a lightweight transformer-enhanced TransUNet for precise lung segmentation, achieving a Dice score of 95.68% on the ”Chest X-ray Masks and Labels” dataset with fewer parameters than traditional transformers. For classification, we employ pre-trained ResNet models (ResNet-50 and ResNet-101) to extract multi-scale feature maps, which are then processed through a modified transformer module to enhance pneumonia detection. This integration of multi-scale feature extraction and lightweight transformer modules ensures robust performance, making our method suitable for resource-constrained clinical environments. Our approach achieves 93.75% accuracy on the ”Kermany” dataset and 96.04% accuracy on the ”Cohen” dataset, outperforming existing methods while maintaining computational efficiency. This work demonstrates the potential of multi-scale transformer architectures to improve pneumonia diagnosis, offering a scalable and accurate solution to global healthcare challenges. https://github.com/amirrezafateh/Multi-Scale-Transformer-Pneumonia
@@ 1 Introduction
Lung segmentation is a critical preprocessing step for improving the accuracy of pneumonia detection in chest X-rays. However, this task faces several challenges. First, the presence of artifacts, overlapping anatomical structures, and low contrast in chest X-rays can make it difficult to accurately delineate lung boundaries sharifi2024da ; agnes2024wavelet . Traditional segmentation models like U-Net, while effective, often struggle to capture fine-grained details and contextual information, leading to suboptimal performance in complex cases yang2024novel . Additionally, the high variability in lung shapes and sizes across patients further complicates the segmentation process. Although transformer-based models have shown promise in addressing these issues, they often come with high computational costs and large parameter counts, making them unsuitable for resource-constrained environments narmadha2025het .
@@
Classification of pneumonia from chest X-rays also presents significant challenges. First, the subtle visual indicators of pneumonia, such as small opacities or localized consolidations, can be easily missed by traditional CNN-based models like ResNet and DenseNet, which primarily focus on local features amin2024automated . While transformer-based models excel at capturing global contextual information, they require large amounts of labeled data and extensive computational resources, limiting their practicality in real-world clinical settings rezvani2025fusionlungnet . Moreover, the domain shift between pre-trained models (e.g., on ImageNet) and medical imaging datasets often results in suboptimal performance, necessitating advanced techniques like transfer learning and domain adaptation chen2017rethinking . These challenges highlight the need for a computationally efficient and accurate model that can effectively leverage both local and global features for pneumonia diagnosis.
@@
To address these challenges, we propose an innovative approach that leverages deep learning with an integrated lightweight transformer, significantly reducing the number of parameters compared to traditional transformers while maintaining lower model complexity. Our method begins with lung segmentation using a TransUNet model, which integrates transformer-based attention mechanisms into the U-Net architecture. The TransUNet model is trained on the ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. Once trained, this pre-trained model is used with frozen weights to predict lung masks for our target datasets, ”Kermany” kermany2018labeled and ”Cohen” cohen2020covid . This segmentation step isolates the lung regions, enhancing the subsequent classification task.
@@
For classification, we utilize pre-trained ResNet models, specifically ResNet-50 and ResNet-101, as the foundation for feature extraction. By drawing multi-scale feature maps from various stages of the ResNet models, we can leverage multiple feature spaces, which enhances the accuracy of our detection. This is achieved through a customized transformer module that employs a cross-attention mechanism, allowing us to make decisions based on more than one feature space. This transformer has been optimized to minimize the number of parameters while preserving performance. By concentrating on the relevant lung regions and integrating multi-scale information, our approach strives to achieve high diagnostic accuracy for pneumonia detection. This architecture reduces the computational load and ensures robust and reliable performance, making it suitable for deployment in resource-limited settings.
@@
  • Reduced number of parameters compared to other state-of-the-art models.
@@ 3 Proposed method
Figure 1: The overview of proposed method
@@ 3.1 Overview
Our classification approach extracts multi-scale feature maps from three key stages of the ResNet models: the outputs of Block 2, Block 3, and Block 4. These stages provide a rich set of features at different scales, which are crucial for accurately identifying Pneumonia in chest X-rays. The extracted feature maps are then processed through a specialized transformer module that employs an attention mechanism. This transformer enhances the feature representation by allowing the network to focus on the most relevant parts of the image.
After the attention processing, the feature maps are concatenated to form a comprehensive representation of the input image. This combined feature map is subsequently fed into fully connected layers to perform the final classification. The overall architecture is designed to effectively integrate multi-scale information and attention mechanisms, thereby improving the classification accuracy. An overview of the proposed method is illustrated in Figure 1.
@@ 3.2.2 Training the TransUNet Model
3.2.3 Applying the Trained TransUNet to the Cohen Dataset
After successfully training the TransUNet model on the ”Chest Xray Masks and Labels” dataset, we utilize this pre-trained segmentation model to predict lung masks for the ”Cohen” dataset. This step is crucial for enhancing the accuracy of the subsequent classification task by focusing on the lung regions within the X-ray images.
The ”Cohen” dataset, which contains chest X-ray images, requires preprocessing to ensure that our classification model focuses on the most relevant regions. To achieve this, we apply the trained TransUNet model to segment the lung areas from these images. The ”Cohen” dataset is first preprocessed to match the input requirements of the TransUNet model. This involves standardizing the image dimensions and normalizing the pixel values to ensure consistency with the training data used for the TransUNet model.
Using the pre-trained TransUNet model, we generate lung masks for each X-ray image in the ”Cohen” dataset. The segmentation model outputs binary masks that highlight the lung regions while suppressing the background. By utilizing the pre-trained TransUNet model to segment the lung regions in the ”Cohen” dataset, we effectively preprocess the data to improve the performance of our classification model. This segmentation step filters out noise and irrelevant features, allowing the classifier to concentrate on the lung areas, thereby enhancing the overall accuracy and robustness of our proposed method.
@@ 3.2.3 Applying the Trained TransUNet to the Cohen Dataset
Following the segmentation of lung regions using the TransUNet model, the next step in our proposed method is the classification task. This task involves accurately identifying the presence of pneumonia or covid19 in the preprocessed chest X-ray images. By focusing on the lung regions isolated during the segmentation phase, we enhance the classification model’s ability to detect relevant features indicative of pneumonia or covid19, thereby improving diagnostic accuracy.
@@ 3.3.2 Transformer
In our proposed method, a transformer is employed to enhance the feature representation obtained from the ResNet backbone, and a similar transformer architecture is used within the TransUNet model. For the classification task, we leverage this transformer to refine the multi-scale feature maps extracted from the ResNet backbone. We begin with two feature maps of size (64,64,64) derived from the ResNet backbone: $B^{\prime 4}$ and $B^{merged}$. Each of these feature maps is fed into a separate transformer, although the structure of the transformers is identical. For simplicity, we will describe the process for $B^{\prime 4}$.
@@
The feature map $B^{\prime 4}$ is reshaped to form the key and value inputs for the transformer. Specifically, as shown in equation 6, $B^{\prime 4}$ is reshaped from $(c,h,w)$ to $(c,h\times w)$.
$B_{flat}^{{}^{\prime}4}=Reshape(B^{\prime 4})$ (6)
@@
Key($K$) and Value($V$): Both the key and value are derived from the reshaped feature map $B_{flat}^{{}^{\prime}4}$
@@ 3.3.3 Output Feature Vector
The output of the transformer for $B^{\prime 4}$ is a feature vector $F^{4}$ of size equal to the number of channels (64 in this case).
The same process is applied to $B^{merged}$, resulting in another feature vector $F^{merged}$ of size 64. The entire process is visually represented in Figure 3, providing a detailed overview of the transformer’s operation on the feature maps.
@@ 4.2 Experimental Setting
We implemented our proposed method using PyTorch version 1.8.1. In classification task, we utilized pre-trained ResNet-50 and ResNet-101 models, which were kept frozen during training to preserve their learned representations. Notably, the learnable parameters of our method amount to only 0.27 million. The model was trained for 30 epochs, and this process was repeated five times to ensure the robustness of the results. The average of these results was then reported.
The Adam optimizer was employed for training, with a learning rate set to $10^{-5}$. All input images were resized to 512 × 512 pixels, and the batch size was set to 64. Notably, no data augmentation techniques were applied; the results reported here were achieved without any such enhancements. All experiments were conducted on an NVIDIA RTX 4090 GPU.
@@ 4.3 Evaluation Metrics
Accuracy=TP+TNTP + TN + FP + $\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP + TN + FP + FN}}$ (10)
@@
[…]
@@ 4.4 Comparison with State-of-the-Art
Proposed method (ResNet-50 as backbone) 96.04%$\pm$0.5 96.70%$\pm$0.5 94.90%$\pm$0.6 95.77%$\pm$0.5
Proposed method (ResNet-101 as backbone) 95.19%$\pm$0.6 95.59%$\pm$0.5 94.19%$\pm$0.6 94.86%$\pm$0.6
@@
Proposed method (ResNet-50 as backbone) 91.67%$\pm$0.6 92.04%$\pm$0.5 94.87%$\pm$0.5 93.43%$\pm$0.6
Proposed method (ResNet-101 as backbone) 93.75%$\pm$0.5 93.98%$\pm$0.5 96.16%$\pm$0.5 95.05%$\pm$0.5
@@ 4.4.2 Classification
The superior performance of our method can be attributed to several key design choices. First, by leveraging pre-trained ResNet models, we extract robust multi-scale feature representations that capture both low-level and high-level patterns in chest X-rays. Freezing the pre-trained layers allows the model to focus on learning task-specific features in the newly added layers, reducing overfitting and improving generalization. Second, our lightweight transformer module enhances the model’s ability to capture global contextual information, which is critical for distinguishing subtle pneumonia-related patterns from complex backgrounds. This integration of local and global feature extraction ensures robust performance across diverse datasets.
@@ 4.5 Ablation study
Transformer Accuracy Precision Recall F1-score

### v5

@@ Abstract
Pneumonia, a prevalent respiratory infection, remains a leading cause of morbidity and mortality worldwide, particularly among vulnerable populations. Chest X-rays serve as a primary tool for pneumonia detection; however, variations in imaging conditions and subtle visual indicators complicate consistent interpretation. Automated tools can enhance traditional methods by improving diagnostic reliability and supporting clinical decision-making. In this study, we propose a novel multi-scale transformer approach for pneumonia detection that integrates lung segmentation and classification into a unified framework. Our method introduces a lightweight transformer-enhanced TransUNet for precise lung segmentation, achieving a Dice score of 95.68% on the ”Chest X-ray Masks and Labels” dataset with fewer parameters than traditional transformers. For classification, we employ pre-trained ResNet models (ResNet-50 and ResNet-101) to extract multi-scale feature maps, which are then processed through Convolutional Residual Attention Module and modified transformer module to enhance pneumonia detection. This integration of multi-scale feature extraction and lightweight attention mechanisms ensures robust performance, making our method suitable for resource-constrained clinical environments. Our approach achieves 93.75% accuracy on the ”Kermany” dataset and 96.04% accuracy on the ”Cohen” dataset, outperforming existing methods while maintaining computational efficiency. https://github.com/amirrezafateh/Multi-Scale-Transformer-Pneumonia
@@ 1 Introduction
Lung segmentation is a crucial preprocessing step for improving the accuracy of pneumonia detection in chest X-rays. However, this task faces several challenges. First, the presence of artifacts, overlapping anatomical structures, and low contrast in chest X-rays can make it difficult to accurately delineate lung boundaries sharifi2024da ; wollek2024out . Traditional segmentation models, such as U-Net, while effective, often struggle to capture fine-grained details and contextual information, resulting in suboptimal performance in complex cases shi2024xraywizard . Additionally, the high variability in lung shapes and sizes across patients further complicates the segmentation process. Although transformer-based models have shown promise in addressing these issues, they often come with high computational costs and large parameter counts, making them unsuitable for resource-constrained environments xie2025efficient ; qiu2025joint .
@@
The classification of pneumonia from chest X-rays also presents significant challenges. First, subtle visual indicators of pneumonia, such as small opacities or localized consolidations, can be easily missed by traditional CNN-based models like ResNet and DenseNet, which primarily focus on local features. While transformer-based models excel at capturing global contextual information, they require large amounts of labeled data and extensive computational resources, limiting their practicality in real-world clinical settings rezvani2025fusionlungnet . Moreover, the domain shift between pre-trained models (e.g., those trained on ImageNet) and medical imaging datasets often results in suboptimal performance, necessitating advanced techniques like transfer learning and domain adaptation chen2017rethinking . These challenges highlight the need for a computationally efficient and accurate model that can effectively leverage both local and global features for pneumonia diagnosis.
@@
To address these challenges, we propose an innovative approach that leverages deep learning through an integrated lightweight transformer, significantly reducing the number of parameters compared to traditional transformers while maintaining lower model complexity. Our method begins with lung segmentation using a TransUNet model, which integrates transformer-based attention mechanisms into the U-Net architecture. The TransUNet model is trained on the ”Chest X-ray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. Once trained, this pre-trained model is used with frozen weights to predict lung masks for our target datasets, ”Kermany” kermany2018labeled and ”Cohen” cohen2020covid . This segmentation step isolates the lung regions, thereby enhancing the subsequent classification task.
@@
For classification, we utilize pre-trained ResNet models, specifically ResNet-50 and ResNet-101, as the foundation for feature extraction. By extracting multi-scale feature maps from various stages of the ResNet models, we can leverage multiple feature spaces, which enhances the accuracy of our detection. This is achieved through a customized transformer module that employs a cross-attention mechanism, allowing us to make decisions based on more than one feature space. This transformer has been optimized to minimize the number of parameters while preserving performance. By concentrating on the relevant lung regions and integrating multi-scale information, our approach aims to achieve high diagnostic accuracy for pneumonia detection. This architecture reduces the computational load and ensures robust and reliable performance, making it suitable for deployment in resource-limited settings.
@@
  • Introducing a Convolutional Residual Attention Module (CRAM) that enriches feature representation by integrating multi-layer residual learning with lightweight attention mechanisms.
@@ 3 Proposed method
Figure 2: The TransUnet Architecture
@@ 3.1 Overview
Our classification approach extracts multi-scale feature maps from three key stages of the ResNet models: the outputs of Block 2, Block 3, and Block 4. These stages provide a rich set of features at different scales, which are crucial for accurately identifying Pneumonia in chest X-rays. The extracted features are first refined through CRAM that enhance discriminative features through dual attention mechanisms and residual learning. These enhanced feature maps are then processed through a specialized transformer module that employs an attention mechanism, further refining the representation by allowing the network to focus on the most relevant parts of the image.
After the attention processing, the feature maps are concatenated to form a comprehensive representation of the input image. This combined feature map is subsequently fed into fully connected layers to perform the final classification. The overall architecture is designed to effectively integrate multi-scale information and attention mechanisms, thereby improving the classification accuracy. The block diagram of the proposed method is illustrated in Figure 1.
@@ 3.2.2 Training the TransUNet Model
Figure 3: Applying trained TransUNet to predict lung masks on ”Cohen”/”Kermay” datasets
@@
3.2.3 Applying the Trained TransUNet to Cohen/Kermany Datasets
As shown in Figure 3, after successfully training the TransUNet model on the ”Chest Xray Masks and Labels” dataset, we utilize this pre-trained segmentation model to predict lung masks for the ”Cohen”/”Kermany” datasets. This step is crucial for enhancing the accuracy of the subsequent classification task by focusing on the lung regions within the X-ray images.
The ”Cohen”/”Kermany” datasets, which contain chest X-ray images, requires preprocessing to ensure that our classification model focuses on the most relevant regions. To achieve this, we apply the trained TransUNet model to segment the lung areas from these images. The ”Cohen” dataset is first preprocessed to match the input requirements of the TransUNet model. This involves standardizing the image dimensions and normalizing the pixel values to ensure consistency with the training data used for the TransUNet model.
Using the pre-trained TransUNet model, we generate lung masks for each X-ray image in the ”Cohen”/”Kermany” datasets. The segmentation model outputs binary masks that highlight the lung regions while suppressing the background. By utilizing the pre-trained TransUNet model to segment the lung regions in the ”Cohen” dataset, we effectively preprocess the data to improve the performance of our classification model. This segmentation step filters out noise and irrelevant features, allowing the classifier to concentrate on the lung areas, thereby enhancing the overall accuracy and robustness of our proposed method.
@@ 3.2.3 Applying the Trained TransUNet to Cohen/Kermany Datasets
Following the segmentation of lung regions using the TransUNet model, the next step in our proposed method is the classification task. This task involves accurately identifying the presence of pneumonia or covid19 in the preprocessed chest X-ray images. By focusing on the lung regions isolated during the segmentation phase, we enhance the classification model’s ability to detect relevant features indicative of pneumonia or covid19, thereby improving diagnostic accuracy. The overview of the proposed method on classification task is shown in Figure 4.
Figure 4: The overview of the proposed method on classification task
@@ 3.3.1 Backbone
Figure 5: The overview of CRAM
3.3.2 Convolutional Residual Attention Module (CRAM)
To further refine the multi-scale feature representations extracted from the ResNet backbone, we propose CRAM. This module operates on two parallel pathways to enhance feature discriminability while preserving spatial relationships. The CRAM is applied to both feature streams: the high-level semantic features from Block 4 ($B^{\prime 4}$) and the merged multi-scale features from Blocks 2 and 3 ($B^{merged}$).
As shown in Figure 5, the CRAM architecture consists of two complementary components working together. For an input feature map $X\in\mathbb{R}^{C\times H\times W}$, the module processes it through both pathways simultaneously.
1. Dual-Attention Mechanism: This module incorporates a sophisticated attention mechanism that operates through two distinct dimensions.
Channel Attention: This component generates a channel attention vector denoted as $F_{CA}$ that emphasizes informative feature channels. The computation is defined as Equation 5. We also define a Sequential Convolution (SC) which have been used on both GAP and GMP that showed in Equation 6.
$F_{CA}(x)=\sigma(GAP(SC)+GMP(SC(x)))\times x$ (5)
$SC=conv_{1\times 1}(Relu(conv_{1\times 1}(x)))$ (6)
where GAP and GMP represent global average pooling and global max pooling operations respectively, SC denotes two-layer 1×1 convolutional network with ReLU activation that reduces channels to $\frac{1}{16}$ and then expands back to original channels. $\sigma$ is the sigmoid function. Finally the channel-refined features are obtained through element-wise multiplication.
Spatial Attention: This component focuses on identifying spatially significant regions within the feature maps. As shown in Equation 7 and Equation 8, it computes both average-pooled and max-pooled features across the channel dimension, concatenates them, and processes them through a convolutional layer with a sigmoid activation to produce spatial attention maps that highlight important spatial locations.
$F_{SA}(x)=\sigma(Relu(conv_{7\times 7}(cat_{SA}))$ (7)
$cat_{SA}=cat(\text{CWA}(F_{CA}(x)),\text{CWM}(F_{CA}(x)))$ (8)
where $F_{CA}$ denotes the output of channel attention part, CWA and CWM represent Channel-Wise Averaging and Channel-Wise Maxing respectively, catcat denotes concatenation.
The final output of the dual-attention mechanism is the element-wise product of the spatially and channel-wise refined features, as formalized in Equation 9.
$\text{DAM}(X)=F_{SA}(x)\times F_{CA}(x)$ (9)
@@ 3.3.2 Convolutional Residual Attention Module (CRAM)
2. Parallel Convolutional Pathway: A complementary convolutional branch, formalized in Equation 10, transforms the input through two successive 3×3 convolutional layers with batch normalization and ReLU non-linearities. This pathway facilitates the learning of complex feature mappings while ensuring stable gradient flow via skip connections.
$\text{Re}(X)=\text{BN}(\text{Conv}_{3\times 3}(\text{ReLU}(\text{BN}(\text{Conv}_{3\times 3}(X)))))$ (10)
where BN denotes Batch Normalization.
The outputs from both pathways are integrated through element-wise addition to produce the final enhanced representation, as formalized in Equation 11. This composite feature enhancement is subsequently applied to both multi-scale feature streams in our architecture, yielding the refined feature maps specified in Equation 12.
$\text{CRAM}(X)=\text{DAM}(X)+\text{Re}(X)$ (11)
$E^{\text{merged}}$ $=\text{CRAM}(B^{\text{merged}})$ (12)
$E^{4}$ $=\text{CRAM}(B^{\prime 4})$
where, $E^{\text{merged}}$ denotes the enhanced multi-scale features from Blocks 2 and 3, while $E^{4}$ represents the refined high-level semantic features from Block 4.
@@ 3.3.3 Transformer
In our proposed method, a transformer is employed to enhance the feature representation obtained from the ResNet backbone, and a similar transformer architecture is used within the TransUNet model. For the classification task, we leverage this transformer to refine the multi-scale feature maps extracted from the ResNet backbone. We begin with two feature maps of size (64,64,64) derived from the ResNet backbone: $E^{4}$ and $E^{merged}$. Each of these feature maps is fed into a separate transformer, although the structure of the transformers is identical. For simplicity, we will describe the process for $E^{4}$.
@@
The feature map $E^{4}$ is reshaped to form the key and value inputs for the transformer. Specifically, as shown in equation 14, $E^{4}$ is reshaped from $(c,h,w)$ to $(c,h\times w)$.
$E_{flat}^{4}=Reshape(E^{4})$ (14)
@@
Key($K$) and Value($V$): Both the key and value are derived from the reshaped feature map $E_{flat}^{4}$
@@ 3.3.4 Output Feature Vector
The output of the transformer for $E^{4}$ is a feature vector $F^{4}$ of size equal to the number of channels (64 in this case).
The same process is applied to $E^{merged}$, resulting in another feature vector $F^{merged}$ of size 64. The entire process is visually represented in Figure 6, providing a detailed overview of the transformer’s operation on the feature maps.
@@ 4.2 Experimental Setting
We implemented our proposed method using PyTorch version 1.8.1. In classification task, we utilized pre-trained ResNet-50 and ResNet-101 models, which were kept frozen during training to preserve their learned representations. Notably, the learnable parameters of our method amount to only 2.29 million. The model was trained for 30 epochs, and this process was repeated five times to ensure the robustness of the results. The average of these results was then reported.
The Adam optimizer was employed for training, with a learning rate set to $10^{-5}$. All input images were resized to 512 × 512 pixels, and the batch size was set to 64. All experiments were conducted on an NVIDIA RTX 4090 GPU.
@@ 4.3 Evaluation Metrics
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP + TN + FP + FN}}$ (18)
@@
$\text{F1 Score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (21)
@@ 4.4 Comparison with State-of-the-Art
[…]
@@
[…]
@@ 4.4.2 Classification
The superior performance of our method can be attributed to several key design choices. First, by leveraging pre-trained ResNet models, we extract robust multi-scale feature representations that capture both low-level and high-level patterns in chest X-rays. Freezing the pre-trained layers allows the model to focus on learning task-specific features in the newly added layers, reducing overfitting and improving generalization. Second, CRAM refine the extracted features through dual attention mechanisms, enhancing discriminative local patterns while suppressing irrelevant information. Our lightweight transformer module then enhances the model’s ability to capture global contextual information, which is critical for distinguishing subtle pneumonia-related patterns from complex backgrounds. This integration of local and global feature extraction ensures robust performance across diverse datasets.
@@ 4.5 Ablation study
CRAM +
Transformer
Accuracy Precision Recall F1-score
@@
Figure 8: Grad-CAM visualizations on Kermany dataset samples
@@
4.6 Explainable AI through Gradient-Weighted Class Activation Mapping
To enhance the interpretability of our model and provide clinical insights into its decision-making process, we employed Gradient-Weighted Class Activation Mapping (Grad-CAM) visualizations. Explainable AI is particularly crucial in medical diagnostics, where understanding the rationale behind a model’s predictions is essential for clinical adoption and trust. Grad-CAM generates heatmaps that highlight the regions of the input image that most significantly influenced the model’s classification decision, effectively creating a visual explanation for the prediction.
As shown in Figure 8, our Grad-CAM analysis reveals the hierarchical feature learning process across different network depths. The first column displays the original input chest X-ray image, providing the baseline for comparison. The subsequent three columns visualize the activation maps from the feature maps extracted at different ResNet blocks: Block 2 (capturing low-level features and basic patterns), Block 3 (capturing mid-level features and structural information), and Block 4 (capturing high-level semantic features and complex patterns). The final two columns present the activation maps from our CRAM: the concatenated feature map (merging Blocks 2 and 3) and the enhanced Block 4 feature map.